A copper calorimeter weighs 100g when empty and 500g when half filled with water at 20oC 50g of dried ice are added and the final temperature of the mixture after all the ice has melted is 10oC. Calculate the specific latent heat of melted ice.


1
Expert's answer
2022-01-03T10:28:56-0500

Solution:


specific heat capacity of water = 4200 "\frac{J}{kgk}" ,

specific heat capacity of copper = 4000 "\frac{J}{kgk}"


latent heat gain by the ice = Heat Lost by water and calorimeter


M(water) = 500g = 0.5kg,

M(calorimeter) = 100g = 0.1kg,

M(ice) = 50g = 0.05kg


mL = Mc@(water) + Mc@(calorimeter)

mL= 0.5*4200*(20-10) + 0.1*4000*(20-10)

mL = 21000 + 4000

mL = 25000

0.05L =25000

L = 500 000 "\frac{J}{kg}"

L (latent heat of water) = 0.5 "\frac{MJ}{kg}"